# Role of intermediate resonances in attosecond photoelectron interferometry in neon

### v1

@@ I Introduction
The reconstruction of attosecond beating by interference of two-photon transitions (RABBIT) technique was demonstrated in 2001 Paul et al. (2001), with the goal to reconstruct the temporal profile of an attosecond pulse train. The oscillations in the photoelectron spectra for the two-color photoionization process as a function of the relative delay $\tau$ between the extreme ultraviolet (XUV) and infrared (IR) pulses contain information about the phase difference between consecutive harmonics, thereby giving access to the attosecond chirp Mairesse et al. (2003); Varjú et al. (2005). This information can be extracted from the experimental data, provided that the target-specific phases imprinted on the photoelectron wave packets are known and can be subtracted from the measured phase differences. In the case of XUV photon energies well above the ionization threshold, and in the absence of resonances in the continuum, these phases are expected to evolve smoothly with the photon energy. In atoms, they can be further decomposed into a photoionization phase, associated with the photoionization process due to the absorption of a single XUV photon, and a continuum-continuum phase due to the interaction of the outgoing photoelectron wave packet with the IR field Dahlström et al. (2012); Dahlström and Lindroth (2014); Pazourek et al. (2015). More recently, the validity of this “decomposition approximation” was generalized and tested also for multiple continuum-continuum transitions Bharti et al. (2021, 2023, 2024).
@@
The presence of intermediate resonances, however, either in the continuum or in the bound-state spectral region, is expected to play a crucial role in attosecond photoelectron interferometry, as these levels can modify the amplitude and phase of one or even both of the two pathways contributing to the interference signal Kheifets and Bray (2021). These effects have been investigated in helium, addressing the role of intermediate bound states reached by single XUV-photon ionization Swoboda et al. (2010); Autuori et al. (2022); Drescher et al. (2022); Neoričić et al. (2022) and of doubly-excited states in the continuum Gruson et al. (2016). In helium, the large energetic distance between consecutive excited levels accessible by single-photon excitation from the ground state, compared to the width of a single harmonic, makes it possible to address individual resonances for each configuration of the harmonics.
@@
In heavier noble-gas atoms on the other hand, the manifold of excited states accessible by dipole transitions from the ground state becomes energetically closer Villeneuve et al. (2017). A single harmonic may thus excite more than one level, thereby generating a rich interference structure in the photoelectron signal in the continuum. Recently, the effect of an intermediate resonance in the interference between two pathways determined by XUV colors was also investigated Carpeggiani et al. (2019); Gryzlova et al. (2022). In molecules, intermediate shape resonances can also affect the phase of the photoemission measured in the laboratory Nandi et al. (2020) and in the recoil frame Heck et al. (2021); Ahmadi et al. (2022) of the target system.
@@
The scheme of the present study is presented in Fig. 1. A comb of XUV harmonics generated by high-order harmonic generation (HHG) in gases includes harmonics below (harmonic $32q-$ 3) and above (harmonics $12q-$ 1 and $12q+$ 1) the ionization threshold of neon. While the harmonic below the ionization threshold can populate a resonant bound excited state, the absorption of a single XUV photon above the ionization threshold leads to the direct emission of a photoelectron (indicated as main lines ML2q-1 and ML2q+1). As a consequence, the first sideband above the ionization threshold (SB2q-2) is formed by the interference of one path influenced by the presence of a manifold of bound states below the ionization threshold and another one resulting from non-resonant ionization above the ionization threshold. We label the process leading to the formation of the sideband $S_{2q-2}$ as “under-threshold” (UT)-RABBIT Kheifets and Bray (2021); Kheifets (2023), because of the involvement of the bound state below the ionization threshold. In contrast, the “above-threshold” (AT)-RABBIT mechanism occurs when both XUV absorption processes already lead to the continuum.
The absorption of a single XUV photon resulting in an electron ejected from the $2p$ subshell of neon leads to the population of continuum states with $s$ and $d$ character, i.e., orbital angular momenta $\ell=0$ and $\ell=2$, respectively Joseph et al. (2020). The subsequent absorption or emission of an IR photon will result in population of final states with $p$ ($\ell=1)$ and $f$ ($\ell=3$) character. The right-hand-side of Fig. 1 shows schematically the character of the different states populated by the absorption of an XUV and IR photon above the ionization threshold of neon. The same conclusions are valid when the first XUV photon populates a bound state below the ionization threshold.
@@ II Experimental Setup
The experimental campaigns were performed using two experimental attosecond sources described in detail in refs. Ahmadi et al. (2020); Ertel et al. (2023). For the first one, the laser driving the HHG was a titanium-doped sapphire system operating at 10 kHz, with a central wavelength of $\lambda=800$ nm delivering pulses with an energy of 1 mJ and a duration (FWHM of intensity profile) of 30 fs. The second one produced IR pulses starting from a Yb:KGW regenerative amplifier with a central wavelength of 1025 nm, 50 kHz repetition rate, and 400 $\mu$J pulse energy. The original pulse duration of 277 fs was reduced to 17.8 fs by exploiting nonlinear broadening in a gas-filled hollow-core fiber (HCF) and subsequent temporal compression using chirped mirrors. Ertel et al. (2023). In the HCF, the type and the pressure of the gas lead to different blue shifts of the central wavelength, thereby offering the possibility to slightly tune the central wavelength in the range between 1001 nm and 1019 nm with an uncertainty of about ±$\pm 5$ nm.
In both beamlines, a delay between the IR pulses and the XUV attosecond pulse train was introduced and controlled with an ultra-stable collinear setup based on drilled glass plates Ahmadi et al. (2020); Ertel et al. (2023).
The XUV and IR pulses were focused by a toroidal mirror in a 3D-momentum-imaging spectrometer (reaction microscope) Ullrich et al. (2003), where they interacted with a supersonic neon gas jet. The products of the ionization were then guided to the photoelectron and photoion detectors by suitable electric and magnetic fields. Here, the peak intensities of the IR pulses at the interaction region of the reaction microscope were estimated as $I_{{\rm IR}}\approx 1\times 10^{12}\,\rm W/cm^{2}$ and $I_{\rm IR}\approx 7\times 10^{11}\,\rm W/cm^{2}$ for the first and second experimental setup, respectively. At the same position, the highest peak intensity of a single attosecond pulse in the train for the two setups was $I_{\rm XUV}\approx 2\times 10^{10}\,\rm W/cm^{2}$ and $I_{\rm XUV}\approx 5\times 10^{9}\,\rm W/cm^{2}$, respectively.
@@
Table 1: Parameters of the polychromatic XUV field. The error values of the amplitudes ${E}_{2q+1}$ are calculated as the standard deviation of the average values around each harmonic maxima. For the phases ${\phi}_{2q+1}$, the error values are evaluated by weighting the standard deviation of the phases around the harmonic maxima with the corresponding amplitude values around the peak.
@@
$S(\theta,\tau)$ $=$ $S_{0}(\theta)+S_{2}(\theta)\cos{(2\,\omega_{\rm IR}\tau-\phi(\theta))}.$
@@
Integrating Eq. (II) over the angles yields
@@ III.1 R-matrix with Time Dependence (RMT)
RMT is an ab-initio technique that solves the time-dependent Schrödinger equation by employing the $R$-matrix paradigm, which divides the configuration space into two separate regions over the radial coordinate of the ejected electron Moore et al. (2011). A general computer code was published Brown et al. (2020), with the most recent version available on GitLab RMT (2023).
In the inner region, close to the nucleus, the time-dependent $N$-electron wave function is represented by a time-independent $R$-matrix (close-coupling) basis with time-dependent coefficients. In the outer region, the wave function is expressed in terms of residual-ion states coupled with the radial wave function of the ejected electron on a finite-difference grid. The solutions in the two separate regions are then matched directly at their mutual boundary. The wave function is propagated in the length gauge of the electric dipole operator. For the typical atomic structure descriptions used in RMT, this converges more quickly than the velocity gauge Hutchinson et al. (2010). On the other hand, propagation in the length gauge usually requires more angular momenta to obtain partial-wave-converged results Cormier and Lambropoulos (1996); Grum-Grzhimailo et al. (2010).
In the nonrelativistic RMT calculations carried out for the present project, the neon target was described within an $R$-matrix inner region of radius 20 (we use atomic units in Sec. III) and an outer region of 15,000. The finite-difference grid spacing in the outer region is 0.08 and the time step for the wave-function propagation was 0.01.
The target was described using the one-electron orbitals generated by Burke and Taylor Burke and Taylor (1975) in their 2-state model, which couples the ionic ground state $(2s^{2}2p^{5})^{2}P$ and the first excited state $(2s2p^{6})^{2}S$. Multiconfiguration expansions, including pseudo-orbitals to account for the term dependence of the physical $2s$ and $2p$ orbitals, were employed for both ionic states as well as the initial $(2s^{2}2p^{6})^{1}S$ state.
@@ III.2 Rate-Equation Approach (RE) and Time-Dependent Perturbation Theory (PT)
Equation (6) may be solved numerically or serves as the starting point for non-stationary perturbation theory: to begin with, one sets $a_{n}^{(0)}(t)=\delta_{n,1}$, which means that initially the system is in the ground state. Then Eq. (6) is integrated to find the first-order expansion coefficients $a_{n}^{(1)}(t)$. Next, these coefficients are put back into Eq. (6) to obtain the second- and even higher-order approximations. While the sideband oscillations observed in this study are determined by the interference of second-order amplitudes, oscillations in the main line arise due to the interference of first- and third-order amplitudes.
@@
The dipole transition discrete-discrete and discrete-continuum matrix elements were obtained in the $jK$-coupling scheme by employing Zatsarinny’s $B$-spline $R$-matrix (BSR) code Zatsarinny (2006), which can be run in electron-ion collision mode with appropriately modified boundary conditions. In this coupling scheme, the quantum numbers characterizing a state are the total electronic angular momentum $J_{f}$ of the ionic core $(2p^{5})^{2}P_{J_{f}}$, the orbital angular momentum $l$ of the electron in the discrete $nl$ or continuum $\varepsilon l$ part of the spectrum, the intermediate quantum number $K$ being the result of the vector addition $\bm{K}=\bm{J_{f}}+\bm{l}$, and, finally, the total electronic angular momentum of the system $\bm{J}=\bm{K}+\bm{1/2}$, where $1/2$ is the electron spin. The choice of the $jK$-scheme made it possible to carefully take into account transitions from discrete to continuum states with different values of the final ionic $J_{f}$ (split by 0.1 eV) in the low-energy region of the spectrum.
@@
The same spectroscopic model as in Popova et al. (2021) was used: all possible terms of configurations $1s^{2}2s^{2}2p^{5}ns$ ($n=3-6$), $1s^{2}2s^{2}2p^{5}nd$ ($n=3-5$), $1s^{2}2s^{2}2p^{4}3s3p$, $1s^{2}2s^{2}2p^{4}3p4s$, and $1s^{2}2s^{1}2p^{5}3s3d$ were included, and the experimental ionization energies Kramida et al. were used.
@@
The calculation of the continuum-continuum transition matrix elements was carried out under the assumption that any correlations between the photoelectron and electrons of the residual ion can be neglected. After the self-consistent wave function of the $(2p^{5})^{2}P$ ion was found with the MCHF software package Froese Fischer et al. (1997), the wave functions of the continuum electron in the frozen field of this ion were produced. Then, using the method described in Mercouris et al. (1996), the continuum-continuum radial matrix elements $R_{\varepsilon l,\varepsilon^{\prime}l^{\prime}}$ were calculated. The relationship between $R_{\varepsilon l,\varepsilon^{\prime}l^{\prime}}$ and the reduced matrix elements in the $jK$-coupling scheme can be obtained using standard formulas from the quantum theory of angular momentum Balashov et al. (2000).
@@
The results for the $d$- to $f$-wave transitions coincide with Gordon’s formula  Gordon (1929) to an accuracy of $2-3$ percent. Being very asymmetric, these transitions proceed mainly by photon absorption and provide the essential contribution to the low-energy part of the spectrum.
@@ IV.1 Delay-averaged data
The lowest energy sideband for the 800 nm wavelength is located around the ionization threshold extending up to about $0.2$ eV in the continuum. Its signal is stronger than that of the adjacent main line ML15. This observation suggests that the sideband is mostly populated through a resonant pathway from below the ionization threshold with the absorption of a photon of the 13th harmonic. Comparison with the tabulated data Kramida et al. strongly suggests that the resonant state is the $(2p^{5}3d)^{1}P$, which can be populated by single-photon absorption from the ground state.
@@
In order to identify the largest contributions from the excited resonances involved in the below-threshold ionization pathway, we carried out a series of calculations within the PT model (specifically, we only “switched on” one of the considered intermediate states and omitted all continuum-continuum transitions). The calculated oscillator strengths for selected transitions between the ground and excited states are presented in Fig. 3(a). The largest contributions are presented in Fig. 3(b), together with the contribution due to absorption of one photon of the $19^{\rm th}$ harmonic and the subsequent emission of an IR photon from ML19.
@@
The pathways leading to excitation of the intermediate $2p^{5}5s$ and $2p^{5}6s$ resonances are found to play only a secondary role (see white bars in Fig. 3(a)). This is (i) consistent with the propensity rule that the absorption of an XUV photon should preferably excite a state with increased orbital angular momentum Fano (1985); Busto et al. (2019) and (ii) ionization cross sections from $d$-states to an $f$-wave ($d\rightarrow f$) are by an order of magnitude larger than the $s\rightarrow p$ and $d\rightarrow p$ cross sections near the threshold.
@@
Figure 4: Parameters for the delay-averaged photoelectron emission at an IR wavelength of 1006 nm: (a,b): spectrum; (c,d): anisotropy parameter $\beta_{2}$; (e,f): anisotropy parameter $\beta_{4}$. The panels on the left column show the near-threshold region on an extended scale. The red points with error bars represent the experimental data, while the curves correspond to the theoretical predictions from RMT (solid black line), PT (blue dashed line), and RE (blue dotted line). In the gaps on the energy axis in panels (d) and (f), the signal is too low for a meaningful extraction of the parameters. The experimental data for the spectra were scaled to provide a good visual fit to the theoretical predictions.
@@
The delay-integrated values of $\beta_{2}$ and $\beta_{4}$ defined according to Eq. (2) are shown together with the predictions of the various theoretical models in panels (c)-(f) of Fig. 4.
For the AT-RABBIT lines SB20 and SB22, there is good agreement in the retrieved $\beta_{2}$ from the experimental data and all theories (see panel (d)). It is also remarkable that all theories show a significant difference between the calculated $\beta_{2}$ for the main lines ML19 and ML21 and the experimental values. Moreover, within the error bars, the latter are much closer to the single-photon values of $\beta_{2}=-0.5$ for ML19 and $-0.16$ for ML21 Codling et al. (1976). This suggests that the IR intensity, and hence the contribution from third-order terms to the modulation in the main lines, may have been overestimated in the modelling.
In general, the agreement is not very good for $\beta_{4}$ (panels (e) and (f)). We note, however, that the determination of $\beta_{4}$ is less reliable, since its extraction from the experimental data is more challenging than that of $\beta_{2}$.
For the UT-RABBIT sideband SB18 (panel (c)), the extracted parameters vary significantly through the line, because several discrete states affect the results to different extent. Note that a resonance often appears in the angular anisotropy parameters in a different way and possibly even at a slightly different energy than for the integrated spectrum Grum-Grzhimailo et al. (2005). For the line-integrated anisotropy parameters, we found $\beta_{2}=1.2$ (expt.), $\beta_{2}=0.9$ (RMT), $\beta_{2}=0.7$ (RE), and $\beta_{2}=0.6$ (PT).

### v2

@@ I Introduction
The reconstruction of attosecond beating by interference of two-photon transitions (RABBIT) technique was demonstrated in 2001 [1], with the goal to reconstruct the temporal profile of an attosecond pulse train. The oscillations in the photoelectron spectra for the two-color photoionization process as a function of the relative delay $\tau$ between the extreme ultraviolet (XUV) and infrared (IR) pulses contain information about the phase difference between consecutive harmonics, thereby giving access to the attosecond chirp [2, 3]. This information can be extracted from the experimental data, provided that the target-specific phases imprinted on the photoelectron wave packets are known and can be subtracted from the measured phase differences. In the case of XUV photon energies well above the ionization threshold, and in the absence of resonances in the continuum, these phases are expected to evolve smoothly with the photon energy. In atoms, they can be further decomposed into a photoionization phase, associated with the photoionization process due to the absorption of a single XUV photon, and a continuum-continuum phase due to the interaction of the outgoing photoelectron wave packet with the IR field [4, 5, 6]. More recently, the validity of this “decomposition approximation” was generalized and tested also for multiple continuum-continuum transitions [7, 8, 9].
@@
The presence of intermediate resonances, however, either in the continuum or in the bound-state spectral region, is expected to play a crucial role in attosecond photoelectron interferometry, as these levels can modify the amplitude and phase of one or even both of the two pathways contributing to the interference signal [10]. These effects have been investigated in helium, addressing the role of intermediate bound states reached by single XUV-photon ionization [11, 12, 13, 14] and of doubly-excited states in the continuum [15]. In helium, the large energetic distance between consecutive excited levels accessible by single-photon excitation from the ground state, compared to the width of a single harmonic, makes it possible to address individual resonances for each configuration of the harmonics.
@@
In heavier noble-gas atoms on the other hand, the manifold of excited states accessible by dipole transitions from the ground state becomes energetically closer [16]. A single harmonic may thus excite more than one level, thereby generating a rich interference structure in the photoelectron signal in the continuum. Recently, the effect of an intermediate resonance in the interference between two pathways determined by XUV colors was also investigated [17, 18]. In molecules, intermediate shape resonances can also affect the phase of the photoemission measured in the laboratory [19] and in the recoil frame [20, 21] of the target system.
@@
The scheme of the present study is presented in Fig. 1. A comb of XUV harmonics generated by high-order harmonic generation (HHG) in gases includes harmonics below (harmonic $32q-$ 3) and above (harmonics $12q-$ 1 and $12q+$ 1) the ionization threshold of neon. While the harmonic below the ionization threshold can populate a resonant bound excited state, the absorption of a single XUV photon above the ionization threshold leads to the direct emission of a photoelectron (indicated as main lines ML2q-1 and ML2q+1). As a consequence, the first sideband above the ionization threshold (SB2q-2) is formed by the interference of one path influenced by the presence of a manifold of bound states below the ionization threshold and another one resulting from non-resonant ionization above the ionization threshold. We label the process leading to the formation of the sideband $S_{2q-2}$ as “under-threshold” (UT)-RABBIT [10, 22], because of the involvement of the bound state below the ionization threshold. In contrast, the “above-threshold” (AT)-RABBIT mechanism occurs when both XUV absorption processes already lead to the continuum.
The absorption of a single XUV photon resulting in an electron ejected from the $2p$ subshell of neon leads to the population of continuum states with $s$ and $d$ character, i.e., orbital angular momenta $\ell=0$ and $\ell=2$, respectively [23]. The subsequent absorption or emission of an IR photon will result in population of final states with $p$ ($\ell=1)$ and $f$ ($\ell=3$) character. The right-hand-side of Fig. 1 shows schematically the character of the different states populated by the absorption of an XUV and IR photon above the ionization threshold of neon. The same conclusions are valid when the first XUV photon populates a bound state below the ionization threshold.
@@ II Experimental Setup
The experimental campaigns were performed using two experimental attosecond sources described in detail in refs. [24, 25]. For the first one, the laser driving the HHG was a titanium-doped sapphire system operating at 10 kHz, with a central wavelength of $\lambda=800$ nm delivering pulses with an energy of 1 mJ and a duration (FWHM of intensity profile) of 30 fs. The second one produced IR pulses starting from a Yb:KGW regenerative amplifier with a central wavelength of 1025 nm, 50 kHz repetition rate, and 400 $\mu$J pulse energy. The original pulse duration of 277 fs was reduced to 17.8 fs by exploiting nonlinear broadening in a gas-filled hollow-core fiber (HCF) and subsequent temporal compression using chirped mirrors. [25]. In the HCF, the type and the pressure of the gas lead to different blue shifts of the central wavelength, thereby offering the possibility to slightly tune the central wavelength in the range between 1001 nm and 1019 nm with an uncertainty of about ±$\pm 5$ nm. In both beamlines, a delay between the IR pulses and the XUV attosecond pulse train was introduced and controlled with an ultra-stable collinear setup based on drilled glass plates [24, 25].
The XUV and IR pulses were focused by a toroidal mirror in a 3D-momentum-imaging spectrometer (reaction microscope) [26], where they interacted with a supersonic neon gas jet. The products of the ionization were then guided to the photoelectron and photoion detectors by suitable electric and magnetic fields. Here, the peak intensities of the IR pulses at the interaction region of the reaction microscope were estimated as $I_{{\rm IR}}\approx 1\times 10^{12}\,\rm W/cm^{2}$ and $I_{\rm IR}\approx 7\times 10^{11}\,\rm W/cm^{2}$ for the first and second experimental setup, respectively. At the same position, the highest peak intensity of a single attosecond pulse in the train for the two setups was $I_{\rm XUV}\approx 2\times 10^{10}\,\rm W/cm^{2}$ and $I_{\rm XUV}\approx 5\times 10^{9}\,\rm W/cm^{2}$, respectively.
@@
Table 1: Parameters of the polychromatic XUV field. The error values of the amplitudes ${E}_{2q+1}$ are calculated as the standard deviation of the average values around each harmonic maximum. For the phases ${\phi}_{2q+1}$, the error values are evaluated by weighting the standard deviation of the phases around the harmonic maxima with the corresponding amplitude values around the peak.
@@
$S(\theta,\tau)$ $=$ $S_{0}(\theta)+S_{2}(\theta)\cos{(2\,\omega_{\rm IR}\tau-\phi(\theta))}.$ (1)
@@
Integrating Eq. (1) over the angles yields
@@ III.1 R-matrix with Time Dependence (RMT)
RMT is an ab-initio technique that solves the time-dependent Schrödinger equation by employing the $R$-matrix paradigm, which divides the configuration space into two separate regions over the radial coordinate of the ejected electron [27]. A general computer code was published [28], with the most recent version available on GitLab [29].
In the inner region, close to the nucleus, the time-dependent $N$-electron wave function is represented by a time-independent $R$-matrix (close-coupling) basis with time-dependent coefficients. In the outer region, the wave function is expressed in terms of residual-ion states coupled with the radial wave function of the ejected electron on a finite-difference grid. The solutions in the two separate regions are then matched directly at their mutual boundary. The wave function is propagated in the length gauge of the electric dipole operator. For the typical atomic structure descriptions used in RMT, this converges more quickly than the velocity gauge [30]. On the other hand, propagation in the length gauge usually requires more angular momenta to obtain partial-wave-converged results [31, 32].
In the nonrelativistic RMT calculations carried out for the present project, the neon target was described within an $R$-matrix inner region of radius 20 (we use atomic units in Sec. III) and an outer region of 15,000. The finite-difference grid spacing in the outer region was 0.08 and the time step for the wave-function propagation was 0.01.
The target was described using the one-electron orbitals generated by Burke and Taylor [33] in their 2-state model, which couples the ionic ground state $(2s^{2}2p^{5})^{2}P$ and the first excited state $(2s2p^{6})^{2}S$. Multiconfiguration expansions, including pseudo-orbitals to account for the term dependence of the physical $2s$ and $2p$ orbitals, were employed for both ionic states as well as the initial $(2s^{2}2p^{6})^{1}S$ state.
@@ III.2 Rate-Equation Approach (RE) and Time-Dependent Perturbation Theory (PT)
Equation (6) may be directly solved numerically and, in the following, we indicate this approach as Rate Equations (RE) for the amplitudes method. Equation (6) can also serve as the starting point for non-stationary perturbation theory: to begin with, one sets $a_{n}^{(0)}(t)=\delta_{n,1}$, which means that initially the system is in the ground state. Then Eq. (6) is integrated to find the first-order expansion coefficients $a_{n}^{(1)}(t)$. Next, these coefficients are put back into Eq. (6) to obtain the second- and even higher-order approximations. In the following, we refer to this second approach as perturbation theory (PT). In PT, the solution of Eq. (6) is expanded into a series as:
$a_{n}(t)=a_{n}^{(0)}+a_{n}^{(1)}+a_{n}^{(2)}….$ (7)
Even- and odd-rank $a_{n}^{(j)}$ have maxima at different photoelectron energies corresponding to SBs (even $j$) and MLs (odd $j$). While the sideband oscillations observed in this study are determined by the interference of second-order amplitudes, oscillations in the mainlines arise due to the interference of first- and third-order amplitudes. The probability for finding an electron in any particular state is defined by the square of the absolute amplitudes. Note that the square of $a_{n}^{(2)}$ and the interference product of $a_{n}^{(1)}$ and $a_{n}^{(3)}$ are of the same order.
@@
The dipole transition discrete-discrete and discrete-continuum matrix elements were obtained in the $jK$-coupling scheme by employing Zatsarinny’s $B$-spline $R$-matrix (BSR) code [34], which can be run in electron-ion collision mode with appropriately modified boundary conditions. In this coupling scheme, the quantum numbers characterizing a state are the total electronic angular momentum $J_{f}$ of the ionic core $(2p^{5})^{2}P_{J_{f}}$, the orbital angular momentum $l$ of the electron in the discrete $nl$ or continuum $\varepsilon l$ part of the spectrum, the intermediate quantum number $K$ being the result of the vector addition $\bm{K}=\bm{J_{f}}+\bm{l}$, and, finally, the total electronic angular momentum of the system $\bm{J}=\bm{K}+\bm{1/2}$, where $1/2$ is the electron spin. The choice of the $jK$-scheme made it possible to carefully take into account transitions from discrete to continuum states with final ionic states that have fine-structure splittings of up to 0.1 eV for different values of $J_{f}$.
@@
The same spectroscopic model as in [35] was used: all possible terms of configurations $1s^{2}2s^{2}2p^{5}ns$ ($n=3-6$), $1s^{2}2s^{2}2p^{5}nd$ ($n=3-5$), $1s^{2}2s^{2}2p^{4}3s3p$, $1s^{2}2s^{2}2p^{4}3p4s$, and $1s^{2}2s^{1}2p^{5}3s3d$ were included, and the experimental ionization energies [36] were used.
@@
The calculation of the continuum-continuum transition matrix elements was carried out under the assumption that any correlations between the photoelectron and electrons of the residual ion can be neglected. After the self-consistent wave function of the $(2p^{5})^{2}P$ ion was found with the MCHF software package [37], the wave functions of the continuum electron in the frozen field of this ion were produced. Then, using the method described in [38], the continuum-continuum radial matrix elements $R_{\varepsilon l,\varepsilon^{\prime}l^{\prime}}$ were calculated. The relationship between $R_{\varepsilon l,\varepsilon^{\prime}l^{\prime}}$ and the reduced matrix elements in the $jK$-coupling scheme can be obtained using standard formulas from the quantum theory of angular momentum [39].
@@
The results for the $d$- to $f$-wave transitions coincide with Gordon’s formula  [40] to an accuracy of $2-3$ percent. Being very asymmetric, these transitions proceed mainly by photon absorption and provide the essential contribution to the low-energy part of the spectrum.
@@ IV.1 Delay-averaged data
The lowest energy sideband for the 800 nm wavelength is located around the ionization threshold extending up to about $0.2$ eV in the continuum. Its signal is stronger than that of the adjacent main line ML15. This observation suggests that the sideband is mostly populated through a resonant pathway from below the ionization threshold with the absorption of a photon of the 13th harmonic. Comparison with the tabulated data [36] strongly suggests that the resonant state is the $(2p^{5}3d)^{1}P$, which can be populated by single-photon absorption from the ground state.
@@
In order to identify the largest contributions from the excited resonances involved in the below-threshold ionization pathway, we carried out a series of calculations within the PT model. Specifically, we only “switched on” one of the considered intermediate states and omitted all continuum-continuum transitions. The calculated oscillator strengths for selected transitions between the ground and excited states are presented in Fig. 3(a). The largest contributions are presented in Fig. 3(b), together with the contribution due to absorption of one photon of the $19^{\rm th}$ harmonic and the subsequent emission of an IR photon from ML19.
@@
The pathways leading to excitation of the intermediate $2p^{5}5s$ and $2p^{5}6s$ resonances are found to play only a secondary role (see white bars in Fig. 3(a)). This is (i) consistent with the propensity rule that the absorption of an XUV photon should preferably excite a state with increased orbital angular momentum [41, 42] and (ii) ionization cross sections from $d$-states to an $f$-wave ($d\rightarrow f$) are by an order of magnitude larger than the $s\rightarrow p$ and $d\rightarrow p$ cross sections near the threshold.
@@
Figure 4: Parameters for the delay-averaged photoelectron emission at an IR wavelength of 1006 nm: (a,b): spectrum; (c,d): anisotropy parameter $\beta_{2}$; (e,f): anisotropy parameter $\beta_{4}$. The panels on the left column show the near-threshold region on an extended scale. The red points with error bars represent the experimental data, while the curves correspond to the theoretical predictions from RMT (solid black line), PT (blue dashed line), and RE (blue dotted line). In the gaps on the energy axis in panels (d) and (f), the signal is too low for a meaningful extraction of the parameters. The experimental data for the spectra were scaled to provide a visual fit to the theoretical predictions.
@@
The delay-integrated values of $\beta_{2}$ and $\beta_{4}$ defined according to Eq. (2) are shown together with the predictions of the various theoretical models in panels (c)-(f) of Fig. 4. For the AT-RABBIT lines SB20 and SB22, there is good agreement in the retrieved $\beta_{2}$ from the experimental data and all theories (see panel (d)). It is also remarkable that all theories show a significant difference between the calculated $\beta_{2}$ for the main lines ML19 and ML21 and the experimental values. Moreover, within the error bars, the latter are much closer to the single-photon values of $\beta_{2}=-0.5$ for ML19 and $-0.16$ for ML21 [43]. This suggests that the IR intensity, and hence the contribution from third-order terms to the modulation in the main lines, may have been overestimated in the modelling. In general, the agreement is not very good for $\beta_{4}$ (panels (e) and (f)). We note, however, that the determination of $\beta_{4}$ is less reliable, since its extraction from the experimental data is more challenging than that of $\beta_{2}$.
For the UT-RABBIT sideband SB18 (panel (c)), the extracted parameters vary significantly through the line, because several discrete states affect the results to different extent. Note that a resonance often appears in the angular anisotropy parameters in a different way and possibly even at a slightly different energy than for the integrated spectrum [44]. For the line-integrated anisotropy parameters, we found $\beta_{2}=1.2$ (expt.), $\beta_{2}=0.9$ (RMT), $\beta_{2}=0.7$ (RE), and $\beta_{2}=0.6$ (PT).